OpenStudy (ujjwal):

An ac circuit of resistor 500 \(\Omega\), an inductor of inductance 0.5H and capacitor of capacitance 8\(\mu\)F in series. The current in the circuit will be maximum when the angular frequency of ac source is: a)500 rad/sec b)\(2\times 10^5\) rad/sec c)4000 Hz d)5000 rad/sec

OpenStudy (anonymous):

Do you know the formula??

OpenStudy (ujjwal):

Unfortunately, I have very less knowledge of ac current!

OpenStudy (kropot72):

Resonance occurs when the reactances of the inductor and capacitor are equal \[\omega L=\frac{1}{\omega C}\] \[\omega ^{2}=\frac{1}{LC}\] \[\omega=\frac{1}{\sqrt{LC}}rad/\sec\]

OpenStudy (ujjwal):

thanks @kropot72

OpenStudy (kropot72):

You're welcome :)